• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Holmstrom

    Log entry time 00:02:25 on July28,2003

    Entry number 107564

    keyword=Swing Shift Summary

    16:00 Start of Shift We are in the middle of Kin. 6.16

    Kin. 6.16
    Right arm momentum = 1.985, angle=12.5 deg.
    Left arm momentum = 3.17, angle=18 deg.
    Septum Current = 184 A

    Raster on
    Beam Current = 4.7 amp

    Right arm run
    Time Run Num. Target PS1 PS2 beam lamba/2 target lamba/2 Events
    15:53-16:09 2906 3He(0 deg) 1 1 IN OUT 3M
    16:10-16:25 2907 3He(0 deg) 1 1 IN OUT 3M
    16:26-16:44 2908 3He(0 deg) 1 1 IN OUT 3M
    16:45-17:01 2909 3He(0 deg) 1 1 IN OUT 3M
    17:02-17:12 2909 3He(0 deg) 1 1 IN OUT 2M
    17:28-17:44 2910 3He(270 deg) 1 1 IN OUT 3M
    17.45-17:53 2911 3He(270 deg) 1 1 IN OUT 1.5M


    16:18 Called MCC and requested that they adjust beam position.

    17:12 Request that MCC block the beam so that we can rotate to (270 deg).
    17:26 Finish target rotation.
    17:27 Request that MCC return the beam.

    17:55 Request MCC mask beam for Kinimatic Change and target change.
    18:02 Beam masked, start kinimatic switch.
    18:41 Kinimatic switch over, asked for beam return.
    18:43 MCC called to say we will get beam, the slit is open, and they estimate bleedthrough at 3%.

    Kin. 6.18
    Right arm momentum = 1.707, angle=12.5 deg.
    Left arm momentum = 3.17, angle=18 deg.
    Septum Current = 158 A

    Raster on
    Beam Current = 5.0 amp

    Right arm run
    Time Run Num. Target PS1 PS2 beam lamba/2 target lamba/2 Events
    18.48-18:58 2913 Reference Empty 1 1 IN OUT 1.35M
    19.07-19:17 2914 Reference N2 5 5 IN OUT 2M
    19.22-19:36 2915 Reference N2 5 5 IN OUT 3M
    19:59-20:14 2917 3He(0 deg) 1 1 IN OUT 3M
    20:15-20:31 2918 3He(0 deg) 1 1 IN OUT 3M
    20:32-20:46 2919 3He(0 deg) 1 1 IN OUT 3M
    20:47-20:56 2920 3He(0 deg) 1 1 IN OUT 2M
    21:08-21:23 2921 3He(270 deg) 1 1 IN OUT 3M
    21:24-21:31 2922 3He(270 deg) 1 1 IN OUT 1.3M


    18:59 Request MCC reduce us to 4.7 uA for N2 reference cell run.
    19:01 Reference cell at 140 psig.
    19:06 The trigger rate is 20khz, so after consulting with RCC, we have set prescales to 5 on the right HRS.

    19:20 We are finished with N2, and we are waiting for the Compton studies to be done, to move to 3He position.
    19:22 To use beam we are takign a 2nd N2 reference cell run.
    19:42 Change target position to 3He Position.
    20:31 Bleedthrough with 5uA is 2.62%.

    20:57 Request MCC mask the beam, so that we can move to 270 deg. orientation of the 3He target.

    21:32 Request MCC mask the beam, so that we can change kinimatics, move to 0 deg. orientation of the 3He target.

    22:02 NMR measurement. P=40.05%

    22:10 Called MCC and request 10 uA for empty reference cell run.

    Kin. 6.20
    Right arm momentum = 1.466, angle=12.5 deg.
    Left arm momentum = 3.17, angle=18 deg.
    Septum Current = 135.5 A

    Raster on
    Beam Current = ?.? amp

    Right arm run
    Time Run Num. Target PS1 PS2 beam lamba/2 target lamba/2 Events
    22.14-22:23 2923 Reference Empty 1 1 IN OUT 1.4M
    22.26-22:36 2924 Reference N2 6 6 IN OUT 1.8M
    22:49-23:02 2925 3He(0 deg) 1 1 IN OUT 3M
    23:03-23:18 2926 3He(0 deg) 1 1 IN OUT 3M
    23:19-23:28 2927 3He(0 deg) 1 1 IN OUT 1.7M
    23:38-23:?? 2928 3He(270 deg) 1 1 IN OUT ?M


    22:23 Called MCC to request 5 uA for the 141 psig N2 reference cell.
    22:37 Called MCC to request a masking of the target, to move the target to pol. 3He position.
    23:22 MCC gives us a bleedthrough of 2.91% with a slit position of 36.2 and a Hall C current of 93 uA.

    23:28 Called MCC to request a mask of the target, so that we can rotate the holding field from 0 deg to 270 deg.
    23:35 Called MCC to request Beam back.

    23:53 MCC called they have a Vaccum Fault in Hall A? They are investigating.